Find the inverse of the following Matrix using Gauss Jordan method

1 2 3 0

2 4 3 2

3 2 1 3

6 8 7 5


1
Expert's answer
2021-05-07T09:55:14-0400

Det A= 1•(20+28+72−16−30−84) −2•(10+42+54−12−45−42)+3•(20+48+72−24−60−48)−0•(28+72+24−36−84−16)=−10−2•7+3•8=0.

Det A=0, thus the inverse matrix isnt exist.